OpenStudy (anonymous):

PLEASE HELP QUICK simplify: sqrt 20/ sqrt 8

OpenStudy (anonymous):

2sqrt(5)/2sqrt(2) = sqrt 5/ sqrt 2

OpenStudy (anonymous):

with work please

OpenStudy (anonymous):

well, 4*5=20 and sqrt 4 =2 then 2*4=8 and sqrt 4 = 2 so that shows the left hand side, does that help?

OpenStudy (anonymous):

\[\sqrt{20} = \sqrt{4*5}= \sqrt{4}*\sqrt{5}=2*\sqrt{5}\]

OpenStudy (anonymous):

I cannot have a sqrt in the denomanator

OpenStudy (anonymous):

so multiply by the fraction sqrt (denom)/sqrt (denom)

OpenStudy (anonymous):

\[\sqrt{2}/\sqrt{2} \]

OpenStudy (anonymous):

the denominator will then be \[\sqrt{2}*\sqrt{2} which equals \sqrt{2}^2 = 2\]

OpenStudy (anonymous):

So... does no reply and posting a new question mean you've got it and indy25 deserves, I dunno, a medal?